SUBMILLIMETER ROTATIONAL SPECTRA OF FORMALDIMINE

\underline{BRIAN J. DROUIN}, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099; ADAM WALTERS, Centre d’Etude Spatiale des Rayonnements, 9, avenue du Colonel Roche - BP 44346 31028 Toulouse Cedex 4, France.

Formaldimine (CH_{2}NH) is a known astrophysical molecule with detections in giant molecular clouds, star-forming regions, hot cores and tentatively an external galaxy. It is also linked to the process of Tholin formation and hence of interest to pre-biotic chemistry schemes. The ground-state rotational spectrum is a simple asymmetric top and has been previously analyzed up to 110 GHz^{a}. We present an experiment designed to study the Tholin formation process using the submillimeter wavelengths (up to 900 GHz) and discuss the resulting new spectroscopy for formaldimine.

aKirchhoff, W. H., Johnson, D. R., & Lovas, F. J. 1973, Journal of Physical and Chemical Reference Data, 2, 1